Circular No. 7519

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

GRB 001109 L. Amati, F. Frontera, and E. Pian, Istituto Tecnologie e Studio Ra-

diazioni Extraterrestri, CNR, Bologna; P. Giommi and M. Capalbi, BeppoSAX Science Data Center, Rome; E. Costa and G. Gandolfi, Istituto di Astrofisica Spaziale, CNR, Rome; J. J. M. in ’t Zand, Space Research Organization Netherlands; C. Guidorzi, Universitá di Ferrara; S. Granata and A. Paolino, BeppoSAX Science Operation Center, Rome, report: “The 2′.5 error box of GRB 001109 provided by the BeppoSAX Wide Field Cameras

(GCN 879; cf. http://gcn.gsfc.nasa.gov/gcn/gcn3 archive.html) was

observed with the BeppoSAX Narrow Field Instruments from Nov. 10.0762 (16.5 hr after the γ-ray burst) to 10.9847 UT. A previously unknown source lying inside the error box has been detected by the Medium Energy Concentrator Spectrometer. The source, designated 1SAX J1830.1+5517, is located at α = 18^{h}30^{m}07^{s}.8, δ = +55^{o}17′56′′ (equinox 2000.0; error radius 50′′) and has an average count rate of 0.00914 ± 0.0007 count/s in the two MECS units, corresponding to a 2–10-keV flux of (7.1 ± 0.5) × 10^{−13} erg cm^{−2} s^{−1}. The flux decreases by a factor of ∼ 2 during the first 20 000 s of the observation. Due to its fading behavior, 1SAX J1830.1+5517 is likely the x-ray afterglow of GRB 001109. We remark that the position of this source is consistent with that of the previously unknown radio source detected in the GRB 001109 error box by Taylor et al. (GCN 880). A second BeppoSAX NFI observation of the field is in progress.”

SUNGRAZING COMETS Further to IAUC 7514 and 7485, D. Hammer reports his measurements

for several additional Kreutz sungrazing comets found by M. Oates on SOHO website images obtained with the C2 coronagraph. The reduced observations and orbital elements by B. G. Marsden are given on the MPECs cited below.

Comet  1997 UT  α_{2000}  δ_{2000}  MPEC
C/1997 T8 Oct. 8.741  12^{h}48.^{m}6  − 6^{o}44′  2000-V35
C/1997 U2  17.181  13 19.9  −10 07  2000-V36
C/1997 U3  20.232  13 31.0  −11 09  2000-V36
C/1997 U4  21.953  13 38.9  −11 50  2000-V36
C/1997 U5  26.862  13 58.0  −13 41  2000-V36
C/1997 U6  26.862  13 57.9  −13 42  2000-V36
C/1997 U7  27.673  14 00.8  −13 54  2000-V36
2000 November 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green